# AI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions

### v1

@@ 4.1 Biases Mitigating Strategies in Forensics
Dror (2013) focused on strategies to mitigate these biases. He proposed that recognizing the spectrum of biases, not only those that can arise from knowing irrelevant case information, but also biases that emerge from base rate regularities, working “backwards” from the suspect to the evidence, and from the working environment itself, can strengthen forensic science. He suggested measures such as a triage approach and case managers to enhance the quality and effectiveness of the work carried out by forensic examiners.
In the pursuit of objectivity and accuracy in forensic analysis, the Forensic Science Regulator of the United Kingdom Government has proposed and implemented several strategies designed to ensure that the analysis is not influenced by any form of bias (Forensic Science Regulator, 2020). One of these strategies is the adoption of a structured and validated approach for the analysis and evaluation of evidence. Two models that exemplify this approach are the ACE-V model and the CAI model. The ACE-V model, which stands for Analysis, Comparison, Evaluation, and Verification, is a widely accepted method for comparing fingerprints (Reznicek et al., 2010). On the other hand, the CAI model, which stands for Case Assessment and Interpretation, is a framework based on Bayesian thinking (Jackson, 2011). It provides clarity on the role of forensic scientists within the criminal justice process and allows for the design of effective, efficient, and robust examination strategies. Documentation, independent verification, and maintaining ethical standards are other key strategies in the forensic process. Documentation involves taking contemporary notes and recording observations, decisions, and changes, ensuring transparency and enabling a thorough review. Independent verification of critical results by a competent peer, unaware of the original examiner’s judgment, adds an extra layer of scrutiny for accuracy. Furthermore, forensic operators must uphold ethical standards such as honesty, integrity, objectivity, and impartiality, and avoid conflicts of interest and external pressures to ensure unbiased and reliable analysis (Forensic Science Regulator, 2020). Lastly, controlling the flow of information to the analyst is paramount. This is achieved using methods such as “sequential unmasking” or “linear unmasking”, which help to prevent unnecessary information from influencing the analyst’s judgment. In this regard, Dror and Kukucka (2021) introduced Linear Sequential Unmasking–Expanded (LSU-E), a methodology aimed at improving decision-making in forensic science by reducing noise and bias. The approach advocates for an initial analysis of raw data or evidence without any reference material or context, followed by a sequential consideration of other relevant information based on its objectivity, relevance, and potential for bias. LSU-E is applicable to all types of decisions within forensic science, not just comparative ones. Its main objective is to optimize the sequence of information exposure to maximize information utility and minimize cognitive and psychological influences. In addition, LSU-E provides guidelines for determining the optimal sequence for presenting task-relevant information and emphasizes the need for transparent documentation of the influence and role of various pieces of information on the decision-making process.
Camilleri et al. (2019) proposed that while some laboratories have implemented bias mitigating strategies with varying impact on operational efficiency, there has been no systematic assessment of the risk posed by cognitive bias. They suggested using a risk management framework to assess the potential impact of bias on forensic interpretations across multiple disciplines. This approach proved useful in assessing the effectiveness of existing bias mitigating strategies and identifying the latent level of risk posed. Furthermore, the authors affirm that all forensic organizations should seek to implement bias limiting measures that are simple, cost-effective, and do not adversely impact efficiency. Recently, Douglass et al. (2023) provided an examination of cognitive biases in digital forensics, particularly focusing on the potential of video-recorded eyewitness identification procedures for assisting evaluators in assessing eyewitness accuracy. The authors conducted three studies involving a total of 1,630 participants. The results revealed that evaluators, on average, successfully differentiated accurate from inaccurate witnesses based on videos of identification procedures alone. However, this ability was compromised when extraneous incriminating evidence was also provided. The authors found that instructions highlighting the limitations of forensic evidence did not preserve evaluators’ ability to discern accuracy when extraneous incriminating case evidence was provided. Moreover, case information affected other judgments, such as perceptions of the witness’s view.
4.2 Biases Mitigating Strategies in Digital Forensics
Grubl and Lallie (2022) addressed one of the most significant challenges in digital forensics: the precise age estimation of victims in child sexual abuse and exploitation cases. Investigators often need to determine the age of victims by examining images and interpreting the sexual development stages and other human characteristics. However, this process can be negatively impacted by a large forensic backlog, cognitive bias, and the immense psychological stress that this work can entail. In response to these challenges, the authors evaluated existing facial image datasets and proposed a new dataset tailored to the needs of similar digital forensic research contributions. This dataset, which included 327 images of individuals aged 0 to 20, is tested on the Deep EXpectation algorithm pre-trained on the IMDB-WIKI dataset (Grubl and Lallie, 2022). The results for young adolescents aged 10 to 15 and older adolescents/adults aged 16 to 20 are very encouraging, achieving Mean Absolute Errors as low as 1.79.
Karie et al. (2019) also delved into the role of cognitive biases in digital forensics, particularly in the context of combating cybercrimes. The authors acknowledged that the increasing complexity of cyber-attacks and the vast amounts of data, also known as Big Data, that investigators need to sift through to unveil Potential Digital Evidence have made the process of forensic investigation more challenging. In response to these challenges, the authors proposed a generic framework for diverging DL cognitive computing techniques into Cyber Forensics (CF), referred to as the DLCF Framework. Such solutions can range from reducing bias in forensic investigations to challenging what evidence is considered admissible in a court of law or any civil hearing.
5 The Art and Science of Deepfakes
@@
The advent of advanced AI technologies has introduced new challenges in this field. One such challenge is the detection of deepfakes, the research area of which is constantly expanding (as shown in Figure 7). Deepfakes are synthetic media created through generative models based mainly on Generative Adversarial Networks (GANs) and Diffusion Models (DMs) (Goodfellow et al., 2014; Sohl-Dickstein et al., 2015). GANs are composed of a Generator (G) and a Discriminator (D) trained simultaneously through a competitive process. The Generator is trained to capture the data distribution of the training set Ts. The Discriminator is trained to distinguish the images created by G from the set Ts. When G creates images with the same data distribution as Ts, D will no longer be able to solve its task and the training phase can be considered completed. Currently, researchers demonstrated that synthetic images created by DMs (Sohl-Dickstein et al., 2015) are better than those generated by GAN engines in terms of photorealism, as the creation process follows a more accurate and “controlled” flow. The basic idea of DMs is to iteratively add noise to an input random noise vector for synthetic data generation in order to model complex data distributions. Figure 8 and Figure 9 show generic GAN and DM schemes related to the creation of synthetic people’s faces.
Deepfakes can pose significant challenges in distinguishing real images from manipulated ones, thereby complicating the task of digital forensic investigators: in fact, the problem of deepfake detection has been addressed extensively by the scientific community (Masood et al., 2023; Verdoliva, 2020). In this context, preventing cognitive biases in digital forensics becomes crucial to ensure the objectivity and neutrality of judgments.
Figure 8: A standard GAN framework including a Generator (G) that creates data samples from random noise, aiming to mimic the training set’s data distribution; a Discriminator (D) differentiates between real and G-generated data. Both are trained competitively.
Figure 9: A generic Diffusion Model learning a latent variable model that maps the latent space with a fixed Markov chain. The data is corrupted by adding Gaussian noise gradually to get the approximate posterior $q(x_{t}|x_{t-1})$. The aim of training is to learn the inverse process $(p_{\theta}(x_{t-1}|x_{t}))$ so that new data can be generated by running the chain backward. The latent variables $x_{1}$, …, $x_{t}$ have the same dimension as the data $x_{0}$.
@@ 6 The Impostor Bias: How AI Media Triggers Bias and Doubt in Perception
The emergence of Generative AI (GenAI) has brought about a sea change in the multimedia landscape, opening up novel avenues for crafting and modifying content. It is largely attributed to the development of a class of machine learning models known as foundation models (FMs) (Rabowsky, 2023). These models, which include the likes of ChatGPT released in November 2022, marked the beginning of a new era in Artificial Intelligence. Foundation models are distinguished by their powerful applications to GenAI, which involves the use of models to generate new content and transform existing content (Rabowsky, 2023). GenAI models can produce high-quality artistic media for visual arts, concept art, music, fiction, literature, video, and animation; distinguishing the real from the fake is becoming increasingly complex, as in the case of human face recognition and its veracity (Figure 10).
Figure 10: Do these people exist? Credit: Guarnera et al. (2022)
The existence of GenAI and its knowledge can lead to the development of a new type of cognitive bias, which we have identified as the “impostor bias”. This refers to an a priori distrust regarding the veracity of multimedia elements such as videos, photos, and audio, due to the awareness that these can be realistically generated by Artificial Intelligence models. Being subject to this bias, a condition towards which we believe we are increasingly heading in parallel with the increase in realism of GenAI multimedia products, entails and will entail a constant doubt that will also be influential when the victims are, for example, court operators, investigators, or magistrates themselves. In this regard, the technologies previously analyzed regarding deepfake detection will increasingly have to orient themselves not only towards the recognition of manipulated or altered multimedia artifacts, but also those generated from scratch starting from descriptive textual inputs. This is a great challenge on which we will have to work to avoid falling into cognitive traps dictated by prejudices of technological suspicion, risking declaring or considering original multimedia products instead generated by AI, and vice versa. Another risk stemming from GenAI pertains to the counterfeiting of artworks, the sale of forged masterpieces, and the infringement of copyright laws. Indeed, such technologies are now capable of simulating the technique and artistic style of the most famous artists, thereby compromising the ability to correctly discern between real and simulated works (Epstein et al., 2023; Leotta et al., 2023) (Figure 11). Forensic examiners from 21 countries showed limited understanding and appreciation of cognitive bias, with fewer than half supporting blind testing, highlighting the need for procedural reforms to blind them to potentially biasing information (Kukucka et al., 2017).
@@
7 Deepfake Detection Methods
@@
Researchers have demonstrated that generative engines leave traces on synthetic content that can be identified and detected in the frequency domain (Guarnera et al., 2020c; Zhang et al., 2019; Marra et al., 2019; Giudice et al., 2021; Dzanic et al., 2020; Durall et al., 2020) (Figure 12). These traces are characterized by both the network architecture (number and type of layers) and its specific parameters (Yu et al., 2019a). In order to distinguish real data from deepfake, Guarnera et al. (2020a, b) proposed methods based on the Expectation-Maximization (Moon, 1996) algorithm capable of capturing traces defined as the correlation of pixels left by convolutional layers.
Wang et al. (2020) used ResNET-50 to distinguish real and ProGAN-generated images, showing generalization across different GANs. FakeSpotter, proposed by Wang et al. (2021), detects GAN-generated faces by monitoring CNN neuron behaviors. Vision Transformer-based solutions for deepfake detection have also been proposed (Wodajo and Atnafu, 2021; Coccomini et al., 2022; Heo et al., 2023; Wang et al., 2022)Wodajo and Atnafu (2021) combined transformers with a convolutional network to extract patches from detected faces in videos. Several studies (Yu et al., 2019b, 2021; Girish et al., 2021; Asnani et al., 2023; Yu et al., 2020; Guarnera et al., 2022) have explored identifying specific GAN models used in creation (Model Attribution Task). Guarnera et al. (2022) distinguished 100 StyleGAN2 instances using ResNET-18 (He et al., 2016) and metric learning (Liu et al., 2012), demonstrating the method’s effectiveness in deepfake model recognition.
@@
Table 3: A summary of the discussed deepfake detection methods. Legend: AUC = Area Under the Curve; TAR = Transfer learning-based Autoencoder with Residuals; ViT = Vision Transformer.
Experimental results of all these methods show that, in general, all generative models leave unique traces that can solve all previously listed tasks with high accuracy. Therefore, these methods can be used in order to help the general user in countering what we called the impostor bias phenomenon. However, we want to highlight an important element of the previously listed methods and not just deepfake detection. All methods in the literature achieve extremely high results in ”constrained” contexts, that is, with architectures known a priori. In practice, current deepfake detectors fail to generalize with synthetic images generated by novel architectures (different from those used during the training procedure), resulting in a drastic drop in classification performance. Some recent new methods (Dong et al., 2023; Coccomini et al., 2023) published by the scientific community seem to be good starting points in order to achieve generalization.

### v2

@@ 4.1 Strategies for Mitigating Cognitive Bias in Forensic Science
In the field of forensic science, cognitive biases can significantly impact the accuracy and impartiality of examiner decisions. Specifically, Dror (2013) focused on strategies to mitigate confirmation bias, contextual influences (Nakhaeizadeh et al., 2014), and base-rate regularities (Thakur et al., 2021). Dror proposed that recognizing the spectrum of biases, not only those that can arise from knowing irrelevant case information, but also biases that emerge from base rate regularities, working “backwards” from the suspect to the evidence, and from the working environment itself, can strengthen forensic science.
To mitigate these effects, several strategies have been proposed. First, cognitive training programs raise awareness among examiners about potential biases. Second, blind verification procedures, where the second examiner is unaware of the first examiner’s decision, help minimize bias. Third, linear examination processes, starting with evidence analysis before considering the suspect, reduce contextual contamination. Fourth, a triage approach tailors procedures based on case complexity, ensuring that resources are allocated effectively. Lastly, cognitive profiles aid in selecting the best-suited individuals for forensic work, enhancing overall objectivity and performance.
In the pursuit of objectivity and accuracy in forensic analysis, the Forensic Science Regulator of the United Kingdom Government has proposed and implemented several strategies designed to ensure that the analysis is not influenced by any form of bias (Forensic Science Regulator, 2020).
  1. 1.
    Blinding Precautions:
    1. (a)
      Analysts should be shielded from information that is not directly relevant to the analysis.
    2. (b)
      Sequential unmasking can be used, where decisions on suitability are made before comparison with reference samples.
    3. (c)
      Careful records should be kept to ensure the order of disclosure and analysis is transparent.
  2. 2.
    Structured Approach (ACE-V and CAI):
    1. (a)
      ACE-V (Analysis, Comparison, Evaluation, and Verification) provides a structured process for fingerprint comparison (Reznicek et al., 2010).
    2. (b)
      CAI (Case Assessment and Interpretation) uses Bayesian thinking and balances prosecution and defense hypotheses (Jackson, 2011).
    3. (c)
      Both approaches emphasize transparency and avoid post hoc rationalization.
  3. 3.
    Awareness, Training, and Competence Assessment:
    1. (a)
      Practitioners need training on cognitive bias risks and mitigation strategies.
    2. (b)
      Proficiency testing and regular assessment help maintain competence.
  4. 4.
    Avoidance of Reconstructive Effects:
    1. (a)
      Contemporaneous notes or technical records prevent reconstructive bias.
    2. (b)
      Analysts should rely on memory as little as possible.
  5. 5.
    Avoidance of Role Effects:
    1. (a)
      Organizational structures should insulate scientists from potential biasing pressures.
    2. (b)
      Scientists must prioritize their duty to the court over any other obligations.
Controlling the flow of information to analysts is crucial to prevent unnecessary influences on their judgment. Dror and Kukucka (2021) introduced Linear Sequential Unmasking–Expanded (LSU-E), a methodology that reduces noise and bias in forensic decision-making. LSU-E involves initial analysis of raw data without reference material, followed by a sequential consideration of relevant information based on objectivity and relevance. This approach optimizes information presentation to enhance utility and minimize cognitive biases. LSU-E also offers guidelines for documenting the influence of information on the decision-making process, ensuring transparency and accountability. Furthermore, Camilleri et al. (2019) emphasized the need for a systematic assessment of cognitive bias risks in forensic laboratories, proposing a risk management framework. Key mitigation strategies include raising awareness through training, developing guidance documents, and limiting access to task-irrelevant information. Redacting irrelevant details, implementing blind known tests, and conducting independent casefile reviews enhance objectivity and reduce expectation bias. These measures may ensure the integrity of forensic interpretations and minimize the impact of cognitive biases.
5 The Intersection of Generative AI and the Craftsmanship of Deepfakes
Generative Artificial Intelligence (GenAI) is an increasingly popular technology that has significant implications across various fields (Bockting et al., 2023; Stokel-Walker and Van Noorden, 2023). It refers to AI systems that can generate new content, such as text, images, and audio, in response to human prompts. These systems, including deepfakes and AI chatbots like Generative Pre-trained Transformers (e.g., GPT-4), use complex algorithms to produce outputs that are often indistinguishable from content created by humans (Stokel-Walker and Van Noorden, 2023). The technology is advancing rapidly, with each new version adding capabilities that increasingly encroach on human skills; however, the use of these “black box” AI tools can introduce biases and inaccuracies, potentially distorting scientific facts while still sounding authoritative.
@@ 5 The Intersection of Generative AI and the Craftsmanship of Deepfakes
Figure 8: A standard GAN framework. A Generator (G) creates data samples from noise, aiming to mimic the training set. A Discriminator (D) differentiates between real and G-generated data. Training ends when D can’t distinguish G’s images from training samples.
The emergence of these sophisticated AI technologies has brought about fresh challenges in this domain. One such challenge is the detection of deepfakes, the research area of which is constantly expanding (as shown in Figure 7). Deepfakes are synthetic media created through generative models based mainly on Generative Adversarial Networks (GANs) and Diffusion Models (DMs) (Goodfellow et al., 2014). GANs are composed of a Generator (G) and a Discriminator (D) trained simultaneously through a competitive process. The Generator is trained to capture the data distribution of the training set Ts. The Discriminator is trained to distinguish the images created by G from the set Ts. When G creates images with the same data distribution as Ts, D will no longer be able to solve its task and the training phase can be considered completed. Currently, researchers demonstrated that synthetic images created by DMs are better than those generated by GAN engines in terms of photorealism, as the creation process follows a more accurate and “controlled” flow. The basic idea of DMs is to iteratively add noise to an input random noise vector for synthetic data generation in order to model complex data distributions. Figure 8 and Figure 9 show generic GAN and DM schemes related to the creation of synthetic people’s faces.
Figure 9: A Diffusion Model architecture. Training data is corrupted with added Gaussian noise. From this data ($x_{T}$ step), a reverse process is constructed to generate new samples resembling the original ones.
Deepfakes can pose significant challenges in distinguishing real images from manipulated ones, thereby complicating the task of digital forensic investigators: in fact, the problem of deepfake detection has been addressed extensively by the scientific community (Masood et al., 2023; Verdoliva, 2020; Lin et al., 2024). In this context, preventing cognitive biases in digital forensics becomes crucial to ensure the objectivity and neutrality of judgments.
@@ 6 The Impostor Bias: How AI Media Triggers Bias and Doubt in Perception
The emergence of GenAI has brought about a sea change in the multimedia landscape, opening up novel avenues for crafting and modifying content. It is largely attributed to the development of a class of machine learning models known as foundation models (FMs) (Rabowsky, 2023). These models, which include the likes of ChatGPT released in November 2022, marked the beginning of a new era in Artificial Intelligence. Foundation models are distinguished by their powerful applications to GenAI, which involves the use of models to generate new content and transform existing content. GenAI models can produce high-quality artistic media for visual arts, concept art, music, fiction, literature, video, and animation; distinguishing the real from the fake is becoming increasingly complex, as in the case of human face recognition and its veracity (Figure 10).
Figure 10: Figures 8 and 9 illustrate the architectures of generative models that produce these images. The verisimilitude of these faces could potentially lead observers to question the existence of the depicted individuals, thereby giving rise to the Impostor Bias. Credit: Guarnera et al. (2022)
The existence of GenAI and its knowledge can lead to the development of a new type of cognitive bias, which we have identified as the “Impostor Bias”. This is a hypothetical bias with no empirical basis yet, and the term “Impostor” is derived from the “Impostor Syndrome”, a psychological phenomenon in which people doubt their competence and fear being exposed as fraudulent. This bias, however, refers to a different context: distrust of multimedia content generated by Artificial Intelligence. In the context of AI, “Impostor Bias” refers to the tendency to doubt the veracity of multimedia elements such as videos, photos and audios, due to the knowledge that these can be realistically generated by AI models. This bias manifests itself as an a priori distrust, regardless of the quality or context of the multimedia content. The name “Impostor Bias” is appropriate because, just as in Impostor Syndrome, there is a persistent doubt about the veracity of something - in this case, AI-generated media content. Although they may appear authentic and realistic, awareness of the possibility that they are AI-generated “impostors” can lead to doubt and distrust.
The term “bias” is used to describe Impostor Bias because it refers to a systematic and predictable tendency in the way people perceive media content, regardless of the evidence presented. For instance, this is not simply a variation in evaluations due to different levels of knowledge of the evaluator. In other words, even when presented with AI-generated media content that is indistinguishable from the real thing, people with Impostor Bias may still doubt its authenticity due to the knowledge that such content may be AI-generated. This is a bias because it is based on an a priori assumption rather than an objective assessment of the content itself. Furthermore, the term “bias” implies that this tendency can lead to distortions in perception and judgement. For example, “Impostor Bias” may lead people to discard or devalue authentic media content because they perceive it as potentially false or misleading.
@@
Judicial practitioners and investigators may struggle with determining the authenticity of multimedia content, especially as AI-generated media becomes more realistic and descriptive. The risk of falling into cognitive traps or suspecting original content as AI-generated underscores the complexity of this issue. Additionally, GenAI raises concerns about artwork counterfeiting, copyright infringement, and the potential for forged masterpieces to be sold as genuine. Indeed, such technologies are now capable of simulating the technique and artistic style of the most famous artists, thereby compromising the ability to correctly discern between real and simulated works (Epstein et al., 2023; Leotta et al., 2023) (Figure 11). Forensic examiners from 21 countries showed limited understanding and appreciation of cognitive bias, with fewer than half supporting blind testing, highlighting the need for procedural reforms to blind them to potentially biasing information (Kukucka et al., 2017).
7 Generative AI and Deepfake Detection Methods
The mitigation of the potential impacts of the hypothetical “Impostor Bias” could significantly rely on Generative AI and deepfake detection methods. As AI-generated media products become increasingly realistic (Verdoliva, 2020; de Lima-Santos and Ceron, 2021), the prevalence of “Impostor Bias” is expected to rise, posing significant challenges in various sectors, including the justice system. In this context, technologies for deepfake detection become increasingly important. These technologies are evolving to not only recognize manipulated or altered multimedia artifacts but also those generated from scratch using descriptive textual inputs.
Detecting content generated by Generative AI, such as ChatGPT, is a rapidly evolving field of study. Recent research has focused on developing machine learning tools capable of distinguishing between human-written text and machine-generated content (Prillaman, 2023; Nature, 2023; Weber-Wulff et al., 2023; Perkins et al., 2024; Liu et al., 2024). These tools, such as the one described in a study by Perkins et al. (2024), analyze various features of writing style, including variation in sentence lengths, and the frequency of certain words and punctuation marks. However, the efficacy of these detection tools can be significantly reduced when they are confronted with machine-generated content that has been modified using techniques designed to evade detection (Perkins et al., 2024). For instance, the study found that the detectors’ already low accuracy rates (39.5%) showed major reductions in accuracy (17.4%) when faced with manipulated content. Despite these challenges, the development of more accurate and robust GenAI detection methods continues to be a critical area of research, given the increasing prevalence of AI-generated content in various domains (Perkins et al., 2024; Liu et al., 2024). For example, a tool called CheckGPT has been developed, which examines 20 features of writing style to determine whether an academic scientist or ChatGPT wrote a piece of text (Liu et al., 2024). The tool was found to be highly accurate, achieving an average classification accuracy of 98% to 99% for task-specific discipline-specific detectors and the unified detectors.
@@ 7 Generative AI and Deepfake Detection Methods
Furthermore, researchers have demonstrated that generative engines leave traces on synthetic content that can be identified and detected in the frequency domain (Guarnera et al., 2020c; Zhang et al., 2019; Marra et al., 2019; Giudice et al., 2021; Dzanic et al., 2020; Durall et al., 2020) (Figure 12). These traces are characterized by both the network architecture (number and type of layers) and its specific parameters (Yu et al., 2019). In order to distinguish real data from deepfake, Guarnera et al. (2020a, b) proposed methods based on the Expectation-Maximization (Moon, 1996) algorithm capable of capturing traces defined as the correlation of pixels left by convolutional layers.
Wang et al. (2020) used ResNET-50 to distinguish real and ProGAN-generated images, showing generalization across different GANs. FakeSpotter, proposed by Wang et al. (2021), detects GAN-generated faces by monitoring CNN neuron behaviors. Vision Transformer-based solutions for deepfake detection have also been proposed (Wodajo and Atnafu, 2021; Coccomini et al., 2022; Heo et al., 2023; Wang et al., 2022)Wodajo and Atnafu (2021) combined transformers with a convolutional network to extract patches from detected faces in videos. Several studies (Yu et al., 2019, 2021; Girish et al., 2021; Asnani et al., 2023; Yu et al., 2020; Guarnera et al., 2022, 2024) have explored identifying specific GAN models used in creation (Model Attribution Task). Guarnera et al. (2022) distinguished 100 StyleGAN2 instances using ResNET-18 (He et al., 2016) and metric learning (Liu et al., 2012), demonstrating the method’s effectiveness in deepfake model recognition.
@@
Table 2: A summary of the discussed deepfake detection methods.
Experimental results of all these methods show that, in general, all generative models leave unique traces that can solve all previously listed tasks with high accuracy. Therefore, these methods can be used in order to help the general user in countering what we called the Impostor Bias phenomenon. However, we want to highlight an important element of the previously listed methods and not just deepfake detection. All methods in the literature achieve extremely high results in “constrained” contexts, that is, with architectures known a priori. In practice, current deepfake detectors fail to generalize with synthetic images generated by novel architectures (different from those used during the training procedure), resulting in a drastic drop in classification performance. Some recent new methods (Dong et al., 2023; Coccomini et al., 2023) published by the scientific community seem to be good starting points in order to achieve generalization.